OpenStudy (anonymous):

A cylindrical capacitor consists of two long concentric metalcylinders.If there is a charge of k coulombs per meter on the inside of the cylinder of radius R1, and -k coulombs per meter on the outside cylinder of radius R2,find the electric field E between the cylinders (Gauss law).What is E inside the inner cylinder?Outside the outer cylinder? Find either by inspection or direct integration the potential ɸ such that E = -  ɸ for each of the three regions above.In each case E is not affected by adding an arbitrary constant to ɸ. Adjust the additive const to make ɸ a continuous func. of space

OpenStudy (anonymous):

E= -∇ɸ This ques. comes under vector analysis, but how to aproach this ques. Pls explain in detail. its an application oif divergence theorem

OpenStudy (anonymous):

by symmetry, E is a function of radius. now use Gauss's law to integration over a cylinder with radius r: E(r)*2pi r l=Qenc/e0 E(r)=Qenc/(2 pi r l e0) if r>R1 and 0 if r<R1 or r>R2 (since no net charge is enclosed for r<R1 or r>R2)

OpenStudy (anonymous):

(Gauss's law is just the divergence theorem applied to electric fields)

OpenStudy (anonymous):

for R2>r>R1, Qenc=k*l, thus E(r)=k/(2 pi r e0)

OpenStudy (anonymous):

Thank you ...the final answer is

OpenStudy (anonymous):

when r greater than or equel to R2, how does it works?

OpenStudy (anonymous):

when r>=R2, net charge (Qenc)=k*l-k*l=0

OpenStudy (anonymous):

the constant term is there so that the potential is continuous

OpenStudy (anonymous):

Thank you very much...

OpenStudy (anonymous):

glad to have helped :)